### v3

@@ Abstract
The advent of the James Webb Space Telescope has revealed a wealth of new galaxies just a few hundred Myr after the Big Bang. Some of these galaxies exhibit unusual N/O ratios that are difficult to explain with stellar populations today. While Wolf-Rayet stars in multiple-burst populations, very massive or rapidly-rotating primordial stars, general relativistic explosions of metal-enriched supermassive stars, or the precursors of globular clusters can in principle account for the nitrogen excess in the galaxies GN-z11 and CEERS 1019, no known stars or supernovae can explain the far higher N/O ratio of 0.46 in GS 3073 at redshift $z=$ 5.55. Here we show that the extreme nitrogen abundances in GS 3073 can be produced by 1000 - 10,000 $M_{\odot}$Ā primordial (Pop III) stars. We find that these are the only candidates that can account for its large N/O ratios and its C/O and Ne/O ratios. GS 3073 is thus the first conclusive evidence in the fossil abundance record of the existence of supermassive Pop III stars at cosmic Dawn.
@@ 1 Introduction
Recent surveys with the James Webb Space Telescope (JWST; Finkelstein etĀ al., 2024; Eisenstein etĀ al., 2023; Bezanson etĀ al., 2024) have revealed the existence of high-redshift galaxies with large N/O abundance ratios that are difficult to explain with populations of stars that are known today. Multiple-bursts of massive star formation (Kobayashi & Ferrara, 2024; Rizzuti etĀ al., 2025; McClymont etĀ al., 2025), very massive or rapidly-rotating primordial stars (Nandal etĀ al., 2024b; Tsiatsiou etĀ al., 2024; Nandal etĀ al., 2024c), general relativistic explosions of metal-enriched supermassive stars (SMSs; Nagele & Umeda, 2023)), or the precursors of globular clusters (Cameron etĀ al., 2023; Marques-Chaves etĀ al., 2024) can in principle account for the supersolar nitrogen to oxygen ratios in the galaxies GN-z11 (Bunker etĀ al., 2023; Senchyna etĀ al., 2024) and CEERS 1019 (Larson etĀ al., 2023). However, no known population of stars or supernovae can explain the far higher N/O ratio of 0.46 recently found in GS 3073 at redshift $z=$ 5.55 (Ji etĀ al., 2024).
@@ 2 Numerical Method
We modeled the evolution of 1000 - 10,000 $M_{\odot}$Ā Pop III stars in 1000 $M_{\odot}$Ā increments with the Geneva Stellar Evolution Code (GENEC; Eggenberger etĀ al., 2008).
@@ 2.1 GENEC
GENEC is a 1D Lagrangian stellar evolution code that uses the Henyey method (Henyey etĀ al., 1965) to solve the equations of stellar structure. Energy generation and nucleosynthesis is calculated with a 38-isotope reaction network and we use the Helmholz equation of state (EOS; Timmes & Swesty, 2000), which includes contributions from degenerate and non-degenerate relativistic and non-relativistic electrons, electron-positron pair production, and radiation. Internal opacities are derived from the OPAL tables (Rogers & Nayfonov, 2002) and convection is approximated by mixing-length theory.
We initialize the stars as 2 $M_{\odot}$Ā hydrostatic $n\sim 3/2$ polytropes with uniform entropy profiles. They have a homogeneous chemical composition with $X=0.7516$, $Y=0.2484$, and $Z=0$ (Ekstrƶm etĀ al., 2012; Nandal etĀ al., 2024a) and a deuterium mass fraction $X_{2}=5\times 10^{-5}$ (HaemmerlĆ© etĀ al., 2018b; Nandal etĀ al., 2023). They are initially evolved with accretion rates above $2\times 10^{-2}$ $M_{\odot}$Ā yr-1, the critical rate above which the protostars grow as cool, red supergiants (Nandal etĀ al., 2023; Herrington etĀ al., 2023). These rates are later increased as the stars approach the Hayashi limit for the first time to ensure that they reach their target masses by the end of the pre-main sequence or at the onset of core hydrogen burning, after which accretion is terminated. We assume cold accretion, in which the entropy of the accreting material is matched to that of the surface of the star (HaemmerlĆ© etĀ al., 2018b) and is assumed to be radiated away, as in geometrically thin, cold disks (HaemmerlĆ© etĀ al., 2013, 2016). Any excess entropy in the infalling matter is radiated away before reaching the surface of the star (Palla & Stahler, 1992; Hosokawa etĀ al., 2010).
As Pop III stars form from pristine gas, we exclude line-driven winds from our models (Vink etĀ al., 2001). We evolve all ten stars to the end of core oxygen burning and a few to the end of silicon burning. The stars are partitioned into up to 420 zones in mass and are adaptively rezoned over their evolution. The stars were also evolved with adaptive time steps, and a moderate convective overshoot of 0.100 in the pre-main sequence and core hydrogen burning stages but not at later stages. Rotation is not included in these models (HaemmerlĆ© etĀ al., 2018a), but see Nandal etĀ al. (2025) for initial attempts to include rotation in SMS evolution. We apply the first-order post-Newtonian Tolman-Oppenheimer-Volkoff (TOV) correction to the equation of hydrostatic equilibrium throughout the evolution of the stars to accommodate GR effects by modifying Newton’s gravitational constant, $G$ (Fuller etĀ al., 1986; Woods etĀ al., 2017; HaemmerlĆ© etĀ al., 2018b):
@@ 3 Elemental Yields
We show HR and Kippenhahn diagrams for the 7000 $M_{\odot}$Ā star as a fiducial case in FigureĀ 1 and its abundance profiles during He burning and at the end of Si burning in FigureĀ 2. The thin radiative layer between the core and the convective H burning shell in panel (h) of FigureĀ 1 is permeable to products of nuclear burning, which pass into the H shell and fuse with hydrogen or are dispersed throughout the star. 12C from He burning reaches the H shell at 1.22 Myr and fuses into 13N and then 14N. This channel creates the N excess consistent with that observed in GS 3073 because it continues throughout helium burning, the second longest evolutionary stage of the star. N is soon dispersed throughout the star by the numerous convective zones above the H layer, becoming the third most abundant element at the surface as shown in FigureĀ 2. 16O due to alpha captures by 12C in the core, which can also create 14N in the H shell, is also mixed throughout the star, becoming the fourth most abundant element at the surface.
14N and 16O synthesis via the CNO cycle depletes 12C over time. Although it is the fourth most abundant element at the start of helium burning it falls to fifth place below 16O halfway through burning as shown in FigureĀ 2. 14N produced in the H shell also returns to the convective core and forms 22Ne after the onset of He burning. This 22Ne is then mixed throughout the star, becoming the sixth most abundant element at the surface. 16O in the core from the CNO cycle can capture an alpha and form 20Ne, but longer 16O formation times result in less 20Ne being produced than 22Ne. Nuclear ash created in the core after the end of central helium burning never reaches the outer layers of the star because convection timescales are longer than later stages of nuclear burning, which last just a few days. Heavier elements such as 26Mg, 28Si, 44Ti, 48Cr and 56Ni thus remain trapped in the core, as shown in FigureĀ 2. Final yields for all 10 stars and mass losses (as explained below) are shown in Table 1 in the Appendix.
Figure 1: HR (left) and Kippenhahn diagrams (right) for the 7000 $M_{\odot}$Ā star. The entire evolution of the star is shown in (a) and (b), its pre-MS evolution is shown in (c) and (d), its core H burning phase is in (e) and (f), and core He burning to the end of Si burning are in (g) and (h). The black and white lines in the Kippenhahn diagrams are isomass and isothermal lines, respectively, and the red and green dots along the track on the lower left in (a) are the zero-age main sequence (ZAMS) positions of 2 - 120 $M_{\odot}$Ā and 1000 - 10,000 $M_{\odot}$Ā Pop III stars for comparison.
Figure 2: Top: abundance profiles of the 7000 $M_{\odot}$Ā star at the beginning (left) and middle (right) of core He burning. Bottom: abundance profiles of the star at the end of Si burning. Purple and white regions are convective and radiative zones, respectively.
@@ 3.1 Abundance Ratios / Dilution Factors
where the $M$ and $A$ are the masses (in $M_{\odot}$) and mass numbers of the elements, respectively. They are plotted versus mass coordinate for the 7000 $M_{\odot}$Ā star at the end of Si burning in FigureĀ 3, in which the $M$ used in EquationĀ 2 is the total mass of the given element lying above the mass coordinate. As discussed in the Appendix, mass loss from these stars is likely but not well quantified, so we treat it as a free parameter and consider three fiducial values: 10% and 50% of the mass of the star and all the mass above its CO core. Thus, in FigureĀ 3 N/O == 1.175, C/O == -0.48 and O/H ++ 12 is 8.65 if 10% of the mass is ejected. The number ratios rise as the mass coordinate falls to 2050 $M_{\odot}$, the CO boundary, after which they diverge in the core because they depend on the evolutionary stage at which the element was created.
We parametrize enrichment of the galaxy by the star with a dilution factor defined to be the mass of the gas contaminated by ejecta from the star divided by the mass of the ejecta. We show N/O, C/O and Ne/O ratios measured against the O/H ++ 12 ratio for all 10 stars as a function of dilution factor from 1 - 1000 for 10% mass loss, 50% mass loss, and the loss of all mass above the CO core in the left panels of FigureĀ 4. He/H ratios, a prominent observational feature of massive stars, and 12C/13C ratios for all the stars and mass losses are discussed in the Appendix. The green star marks the observed abundance ratio for GS 3073 derived from rest-frame UV emission lines (Ji etĀ al., 2024). The yellow and brown stars are corresponding ratios for CEERS 1019 and GN-z11 (Marques-Chaves etĀ al., 2024; Cameron etĀ al., 2023). Illumination by a Type-1 active galactic nucleus (AGN; Ji etĀ al., 2024) in GS 3073, and its lower redshift, allow these ratios to be tightly constrained.
@@
Figure 4: N/O, C/O and Ne/O number ratios for 1000 - 10,000 $M_{\odot}$Ā Pop III stars. Top: 10% mass loss. Center: 50% mass loss. Bottom: loss of all the mass above the CO core. The green, yellow and brown stars indicate corresponding ratios for GS 3073, CEERS 1019 and GN-z11, respectively. The dilution factor is the mass of the gas contaminated by outflows from the star divided by the mass of the outflow.
@@ 3.2 N/O Ratios
For mass losses of 10% our stars produce N/O ratios higher than that of GS 3073 except for 9000 $M_{\odot}$, which has an N/O ratio of 0.39 at dilution factors of 20 - 30 that falls within the error bars for GS 3073. The second closest model is the 8000 $M_{\odot}$Ā star, with an N/O ratio of 0.60 at dilution factors of 5 - 15. For mass losses of 50%, 8 of the 10 stars produce N/O ratios above those of GS 3073 except for the 9000 and 10,000 $M_{\odot}$Ā models. The N/O ratios of the 8000 and 9000 $M_{\odot}$Ā stars, 0.52 and 0.41, lie within the error bars for GS 3073 at dilution factors of 24 - 46. If the star sheds all its mass above the CO core only the 3000, 5000 and 7000 $M_{\odot}$Ā stars produce N/O ratios greater than those of GS 3073. The others still exhibit relatively high N/O ratios, such as 0.1 for the 9000 $M_{\odot}$Ā model. The 1000 and 3000 $M_{\odot}$Ā stars closely match the observed abundances, with N/O ratios of 0.38 and 0.48 for dilution factors of 12 - 38 and 31 - 56, respectively.
@@ 3.3 C/O Ratios
As shown in the upper right panels of FigureĀ 4, four stars with 10% mass loss produce an overabundance of C to O compared to those of GS 3073. The 2000, 5000 and 7000 $M_{\odot}$Ā stars have C/O ratios that are closest to the observed values, -0.45, -0.44 and -0.14, respectively at dilution factors of 5 - 15 at 2000 $M_{\odot}$Ā but only at a dilution factor of 1 for 5000 and 7000 $M_{\odot}$. For the 50% case, three models produce an overabundance of C/O with respect to GS 3073. The 7000 $M_{\odot}$Ā model most closely matches GS 3073 with a C/O ratio of -0.30 for dilution factors of 1 - 8. For the loss of all mass above the CO core, only 2 stars produce overabundances of C/O, with the 7000 $M_{\odot}$Ā model again most closely matching GS 3073 with a value of -0.28 over dilution factors of 1 - 8.

### v4

@@ Abstract
The advent of the James Webb Space Telescope has revealed a wealth of new galaxies just a few hundred Myr after the Big Bang, a few of which exhibit unusual N/O ratios that are difficult to explain with stellar populations today. While Wolf-Rayet stars in multiple-burst populations, very massive or rapidly-rotating primordial stars, general relativistic explosions of metal-enriched supermassive stars, or the precursors of globular clusters can in principle account for the nitrogen excess in the galaxies GN-z11 and CEERS 1019, no known stars or supernovae can explain the far higher N/O ratio of 0.46 in GS 3073 at redshift $z=$ 5.55. Here we show that the extreme nitrogen abundances in GS 3073 can be produced by 1000 - 10,000 $M_{\odot}$ primordial (Pop III) stars. We find that these are the only candidates that can account for its large N/O ratios and its C/O and Ne/O ratios. GS 3073 is thus the first conclusive evidence in the fossil abundance record of the existence of supermassive Pop III stars at cosmic Dawn.
@@ 1 Introduction
Recent surveys with the James Webb Space Telescope (JWST; Finkelstein et al., 2024; Eisenstein et al., 2023; Bezanson et al., 2024) have revealed the existence of high-redshift galaxies with large N/O abundance ratios that are difficult to explain with populations of stars that are known today. Multiple-bursts of massive star formation (Kobayashi & Ferrara, 2024; Rizzuti et al., 2025; McClymont et al., 2025), very massive or rapidly-rotating primordial stars (Nandal et al., 2024b; Tsiatsiou et al., 2024; Nandal et al., 2024c), general relativistic explosions of metal-enriched supermassive stars (SMSs; Nagele & Umeda, 2023)), or the precursors of globular clusters (Cameron et al., 2023; Marques-Chaves et al., 2024) can in principle account for the supersolar nitrogen to oxygen ratios in the galaxies GN-z11 (Bunker et al., 2023; Senchyna et al., 2024) and CEERS 1019 (Larson et al., 2023). However, no known population of stars or supernovae can explain the far higher N/O ratio of 0.46 recently found in GS 3073 at redshift $z=$ 5.55 (Ji et al., 2024).
@@ 2 Numerical Method
We modeled the evolution of 1000 - 10,000 $M_{\odot}$ Pop III stars in 1000 $M_{\odot}$ increments with the Geneva Stellar Evolution Code (GENEC; Eggenberger et al., 2008).
@@ 2.1 GENEC
GENEC is a 1D Lagrangian stellar evolution code that uses the Henyey method (Henyey et al., 1965) to solve the equations of stellar structure. Energy generation and nucleosynthesis is calculated with a 38-isotope reaction network and we use the Helmholz equation of state (EOS; Timmes & Swesty, 2000), which includes contributions from degenerate and non-degenerate relativistic and non-relativistic electrons, electron-positron pair production, and radiation. Internal opacities are derived from the OPAL tables (Rogers & Nayfonov, 2002) and convection is approximated by mixing-length theory.
We initialize the stars as 2 $M_{\odot}$ hydrostatic $n\sim 3/2$ polytropes with uniform entropy profiles. They have a homogeneous chemical composition with $X=0.7516$, $Y=0.2484$, and $Z=0$ (Ekström et al., 2012; Nandal et al., 2024a) and a deuterium mass fraction $X_{2}=5\times 10^{-5}$ (Haemmerlé et al., 2018b; Nandal et al., 2023). They are initially evolved with accretion rates above $2\times 10^{-2}$ $M_{\odot}$ yr-1, the critical rate above which the protostars grow as cool, red supergiants (Hosokawa et al., 2012, 2013; Nandal et al., 2023; Herrington et al., 2023). These rates are later increased as the stars approach the Hayashi limit for the first time to ensure that they reach their target masses by the end of the pre-main sequence or at the onset of core hydrogen burning, after which accretion is terminated. We assume cold accretion, in which the entropy of the accreting material is matched to that of the surface of the star (Haemmerlé et al., 2018b) and is assumed to be radiated away, as in geometrically thin, cold disks (Haemmerlé et al., 2013, 2016). Any excess entropy in the infalling matter is radiated away before reaching the surface of the star (Palla & Stahler, 1992; Hosokawa et al., 2010). For fixed final mass and fixed ejection rule, the integrated ejecta are insensitive to whether pre–MS accretion is variable or constant. By the end of core–He burning the differences are $\leq 10^{-2}$ in mass fraction for N, O, and Ne, and by core–Si burning they are $\leq 10^{-4}$; the only systematic change is the total age.
As Pop III stars form from pristine gas, we exclude line-driven winds from our models (Vink et al., 2001). We evolve all ten stars to the end of core oxygen burning and a few to the end of silicon burning. The stars are partitioned into up to 420 zones in mass and are adaptively rezoned over their evolution. The stars were also evolved with adaptive time steps, and a moderate convective overshoot of 0.100 in the pre-main sequence and core hydrogen burning stages but not at later stages. Rotation is not included in these models (Haemmerlé et al., 2018a), but see Nandal et al. (2025) for initial attempts to include rotation in SMS evolution. We apply the first-order post-Newtonian Tolman-Oppenheimer-Volkoff (TOV) correction to the equation of hydrostatic equilibrium throughout the evolution of the stars to accommodate GR effects by modifying Newton’s gravitational constant, $G$ (Fuller et al., 1986; Woods et al., 2017; Haemmerlé et al., 2018b):
@@
In contrast to Nandal et al. (2024b), we follow each model at least to core–O burning and for a subset to core–Si burning, which fixes the O and Ne stratification outside the CO core and enables robust Ne/O predictions used below.
@@ 3 Elemental Yields
Figure 1: HR (top) and Kippenhahn diagrams (bottom) for the 7000 $M_{\odot}$ star. The entire evolution of the star is shown in (a) and (b), its pre-MS evolution is shown in (c) and (d), its core H burning phase is in (e) and (f), and core He burning to the end of Si burning are in (g) and (h). The black and white lines in the Kippenhahn diagrams are isomass and isothermal lines, respectively, and the red and green dots along the track on the lower left in (a) are the zero-age main sequence (ZAMS) positions of 2 - 120 $M_{\odot}$ and 1000 - 10,000 $M_{\odot}$ Pop III stars for comparison.
Figure 2: Abundance profiles of the $7000\,M_{\odot}$ star at three evolutionary stages. Top: beginning of core He burning. Middle: middle of core He burning. Bottom: end of Si burning. Purple and yellow regions mark convective and radiative zones, respectively.
We show HR and Kippenhahn diagrams for the 7000 $M_{\odot}$ star as a fiducial case in Figure 1 and its abundance profiles during He burning and at the end of Si burning in Figure 2. The thin radiative layer between the core and the convective H burning shell in panel (h) of Figure 1 is permeable to products of nuclear burning, which pass into the H shell and fuse with hydrogen or are dispersed throughout the star. 12C from He burning reaches the H shell at 1.22 Myr and fuses into 13N and then 14N. This channel creates the N excess consistent with that observed in GS 3073 because it continues throughout helium burning, the second longest evolutionary stage of the star. N is soon dispersed throughout the star by the numerous convective zones above the H layer, becoming the third most abundant element at the surface as shown in Figure 2. 16O due to alpha captures by 12C in the core, which can also create 14N in the H shell, is also mixed throughout the star, becoming the fourth most abundant element at the surface.
14N and 16O synthesis via the CNO cycle depletes 12C over time. Although it is the fourth most abundant element at the start of helium burning it falls to fifth place below 16O halfway through burning as shown in Figure 2. 14N produced in the H shell also returns to the convective core and forms 22Ne after the onset of He burning. This 22Ne is then mixed throughout the star, becoming the sixth most abundant element at the surface. 16O in the core from the CNO cycle can capture an alpha and form 20Ne, but longer 16O formation times result in less 20Ne being produced than 22Ne. Nuclear ash created in the core after the end of central helium burning never reaches the outer layers of the star because convection timescales are longer than later stages of nuclear burning, which last just a few days. Heavier elements such as 26Mg, 28Si, 44Ti, 48Cr and 56Ni thus remain trapped in the core, as shown in Figure 2. Final yields for all 10 stars and mass losses (as explained below) are shown in Table 1 in the Appendix.
@@ 3.1 Abundance Ratios / Dilution Factors
where the $M$ and $A$ are the masses (in $M_{\odot}$) and mass numbers of the elements, respectively. They are plotted versus mass coordinate for the 7000 $M_{\odot}$ star at the end of Si burning in Figure 3, in which the $M$ used in Equation 2 is the total mass of the given element lying above the mass coordinate. As discussed in the Appendix, mass loss from these stars is likely but not well quantified, so we treat it as a free parameter and consider three fiducial values: 10% and 50% of the mass of the star and all the mass above its CO core. Thus, in Figure 3 N/O == 1.175, C/O == -0.48 and O/H ++ 12 is 8.65 if 10% of the mass is ejected. The number ratios rise as the mass coordinate falls to 2050 $M_{\odot}$, the CO boundary, after which they diverge in the core because they depend on the evolutionary stage at which the element was created.
We parametrize enrichment of the galaxy by the star with a dilution factor defined to be the mass of the gas contaminated by ejecta from the star divided by the mass of the ejecta. We show N/O, C/O and Ne/O ratios measured against the O/H ++ 12 ratio for all 10 stars as a function of dilution factor from 1 - 1000 for 10% mass loss, 50% mass loss, and the loss of all mass above the CO core in the left panels of Figure 4. He/H ratios, a prominent observational feature of massive stars, and 12C/13C ratios for all the stars and mass losses are discussed in the Appendix. The green star marks the observed abundance ratio for GS 3073 derived from rest-frame UV emission lines (Ji et al., 2024). The yellow and brown stars are corresponding ratios for CEERS 1019 and GN-z11 (Marques-Chaves et al., 2024; Cameron et al., 2023). Illumination by a Type-1 active galactic nucleus (AGN; Ji et al., 2024) in GS 3073, and its lower redshift, allow these ratios to be tightly constrained.
Nitrogen enrichment follows a robust sequence in our supermassive models: 12C produced in the convective He core permeates the thin radiative buffer into the convective H–burning shell, where the CNO cycle converts it to 14N; intermediate convective zones above the shell then disperse 14N through the envelope (see Figs. 1 and 2). The appearance of these zones is controlled by the standard convective instability criterion, $\nabla_{\rm rad}>\nabla_{\rm ad}$, with the radiative gradient
$\nabla_{\rm rad}\;=\;\frac{3\,\kappa\,L\,P}{16\,\pi\,a\,c\,G\,m\,T^{4}}\,$,
so that near–Eddington luminosities in SMS envelopes favor extended convection without requiring any ad hoc mixing. After accretion ceases, the structure relaxes on a Kelvin–Helmholtz timescale $\tau_{\rm KH}\!=\!GM^{2}/(RL)$ that is short compared to the duration of core-He burning; as a result the convective topology and the 12C supply converge across $10^{3}$$10^{4}$ M. This produces high N/O with only weak stellar mass dependence in the fixed fraction mass-loss cases (Top and Center panels of Fig. 4). By contrast, in the “above CO core” prescription a small change in the mass cut across the steep N rich/O poor layer just outside the CO core can yield large local excursions in N/O, reflecting a dependence on the composition gradient, rather than on a progenitor mass trend.
@@
Figure 4: N/O, C/O and Ne/O number ratios for 5 select models in the mass range of 1000 - 10,000 $M_{\odot}$ Pop III stars. Top: 10% mass loss. Center: 50% mass loss. Bottom: loss of all the mass above the CO core. The green, yellow and brown stars indicate corresponding ratios for GS 3073, CEERS 1019 and GN-z11, respectively. The dilution factor is the mass of the gas contaminated by outflows from the star divided by the mass of the outflow. Outliers appear only in the “CO core” case (bottom panel), where the mass cut can intersect the narrow N rich/O poor layer immediately outside the core; the 10% and 50% fixed fraction cases average over this gradient and exhibit much smaller scatter.
@@ 3.2 N/O Ratios
For mass losses of 10% our stars produce N/O ratios higher than that of GS 3073 except for 9000 $M_{\odot}$, which has an N/O ratio of 0.39 at dilution factors of 20 - 30 that falls within the error bars for GS 3073. The second closest model is the 8000 $M_{\odot}$ star, with an N/O ratio of 0.60 at dilution factors of 5 - 15. For mass losses of 50%, 8 of the 10 stars produce N/O ratios above those of GS 3073 except for the 9000 and 10,000 $M_{\odot}$ models. The N/O ratios of the 8000 and 9000 $M_{\odot}$ stars, 0.52 and 0.41, lie within the error bars for GS 3073 at dilution factors of 24 - 46. If the star sheds all its mass above the CO core only the 3000, 5000 and 7000 $M_{\odot}$ stars produce N/O ratios greater than those of GS 3073. The others still exhibit relatively high N/O ratios, such as 0.1 for the 9000 $M_{\odot}$ model. The 1000 and 3000 $M_{\odot}$ stars closely match the observed abundances, with N/O ratios of 0.38 and 0.48 for dilution factors of 12 - 38 and 31 - 56, respectively. The extended evolution allows us to match the observed $\log(\mathrm{Ne/O})$ of GS 3073 while preserving the near–vertical rise in N/O at roughly fixed O/H (Ji et al., 2024); this diagnostic was unavailable in our earlier work because the models did not reach advanced burning.
@@ 3.3 C/O Ratios
As shown in the upper right panels of Figure 4, four stars with 10% mass loss produce an overabundance of C to O compared to those of GS 3073. The 2000, 5000 and 7000 $M_{\odot}$ stars have C/O ratios that are closest to the observed values, -0.45, -0.44 and -0.14, respectively at dilution factors of 5 - 15 at 2000 $M_{\odot}$ but only at a dilution factor of 1 for 5000 and 7000 $M_{\odot}$. For the 50% case, three models produce an overabundance of C/O with respect to GS 3073. The 7000 $M_{\odot}$ model most closely matches GS 3073 with a C/O ratio of -0.30 for dilution factors of 1 - 8. For the loss of all mass above the CO core, only 2 stars produce overabundances of C/O, with the 7000 $M_{\odot}$ model again most closely matching GS 3073 with a value of -0.28 over dilution factors of 1 - 8.